A 30 m steel tape was standardised at a temperature of 20o C and under a pull 5 kg. The tape was used in catenary to fix a distance of 28 m between two points at 40o C and under a pull of 5 kg. Given that the cross-sectional area of the tape = 0.02 cm2, total weight 470 g. Young’s modulus of steel = 2.1 X 106 kg/cm, and coefficient of linear expansion = 11 x10-6per oC, (a) find the correct distance between the points, and (b) find the value of pull for which the measured distance would be equal to the correct distance.



1
Expert's answer
2022-02-15T14:32:02-0500

Δl=l1αΔt


l1​="\frac{Δl}{αΔt}" ​l2​−l1​=l1​αΔt


l2​=l1​(1+αΔt)


l1​="\frac{l2}{1+αΔt}" = "\frac{(28\times 11) 10^6}{(2.1 X 10^6)\times20)}"


​I​1="\frac{308,000,000}{42,000,000}" m= 7.33m


=7.33m